UT 20160322 LUCI1 AO Nighttime Commissioning

Observer LUCI-Team: JHeidt, APramskiy, WSeifert, DThompson,
Observer AO: DMiller, EPinna (remote form Italy), APuglisi (remote from Italy), GTaylor
Telescope Operator: DHuerta
LBTO Support: TSargent

Plan for the night

- Observing conditions are expected to vary strongly due to Cirrus and wind, thus:

- Do AFC-tests imaging mode on sky

- Performance of LUCI/FLAO at varying observing conditions

- Some more FLAO-tests for optical gain optimization, if seeing is leq 1"

Summary

- A little bit of AFC and FLAO testing

- 2/3 of the night lost due to wind

Details

All times are UT
Files are luci1.20160322.NNNN.fits

01:35 Sunset, quite a few Cirrus around, T = +6 deg, Hum = 25%, variable wind 7-14 m/s

01:50 David checks pointing and collimates

Stability of system for fast moving objects, accuracy of objects

02:23 Tried to preset to Psyche (16), p16_BgA_0230_22mar2016, failed, restartet telescope manager, tried again, faield again,
preset from a script which was used successfully executed last night, failed too, Alexander needs to check, rollback to earlier version

02:46 Preset to Psyche (16), p16_BgA_0230_22mar2016 (from V2-set of scripts), OK, RunAO, failed,
preset to p16_BgA_0300_22mar2016 (from V2-set of scripts), OK, RunAO failed, RS star was there but disappeared,
preset again, star there, RunAO failed, again, David checked pointing, preset again, didn't work, perhaps
a problem with the GS

  • AL sent 3 different versions of his scripts within one hour, the last set exactly at the time we were ready to start
  • Since we are doing a test not related to commissioning, this was an exception,
  • and will NOT be accepted next time (even the last set of scripts had an obvious error)

Testing Gopt by Greg & Doug

03:11 Doug and Greg take over for optical Gain optimization

Sky full of heavy Cirrus

Test AFC

04:44 Preset to M82, OK, RunAO, OK, "locked " on pinhole for AFC, OK, cont'd, offset failed,
...skip frame... just before the offset (reason was FLAO not LUCI!), OK started, bin=2, 100 modes due to extinction, Ks SRR 180sec, 10samples, #0087-0094

05:54 Need to close due to wind

Com#5 over

-- %USERSIG{JochenHeidt - 2016-03-22}%
Topic revision: r5 - 22 Mar 2016, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback